# Insights from the Inverse: Reconstructing LLM Training Goals Through Inverse Reinforcement Learning

### v2

@@ Abstract
Large language models (LLMs) trained with Reinforcement Learning from Human Feedback (RLHF) have demonstrated remarkable capabilities, but their underlying reward functions and decision-making processes remain opaque. This paper introduces a novel approach to interpreting LLMs by applying inverse reinforcement learning (IRL) to recover their implicit reward functions. We conduct experiments on toxicity-aligned LLMs of varying sizes, extracting reward models that achieve up to 85% accuracy in predicting human preferences. Our analysis reveals key insights into the non-identifiability of reward functions, the relationship between model size and interpretability, and potential pitfalls in the RLHF process. We demonstrate that IRL-derived reward models can be used to fine-tune new LLMs, resulting in comparable or improved performance on toxicity benchmarks. This work provides a new lens for understanding and improving LLM alignment, with implications for the responsible development and deployment of these powerful systems.
@@ 1 Introduction
Inverse Reinforcement Learning (IRL) (Ng & Russell, 2000) is widely used in robotics and control theory to infer latent reward structures from observed behavior. By treating observed actions as expert demonstrations, IRL techniques aim to reverse-engineer the hidden reward function that an agent is implicitly optimizing. In this work, we propose a novel application of IRL to LLMs trained via RLHF. We posit that if the outputs of an RLHF-trained LLM can be interpreted as demonstrations from an “expert” policy, then IRL methods—particularly those based on maximum margin formulations—can be employed to recover the hidden reward functions that guided the training process. This approach provides insights into decision-making processes and enhances model auditing.
@@
We employ Maximum Margin IRL to learn a reward model $\hat{R}:\mathcal{O}\rightarrow\mathbb{R}$ that distinguishes between preferred (non-toxic) and suboptimal (toxic) responses. By parameterizing $\hat{R}$ using hidden representations from a base LLM and optimizing it with an asymmetric margin loss, we enforce clear reward separation while addressing non-identifiability challenges. Our experiments focus on toxicity reduction, evaluating Pythia models (70M and 410M parameters) on Jigsaw Toxicity and RealToxicityPrompts benchmarks. Results show that IRL-extracted reward functions closely align with human judgments. We quantitatively evaluate the recovered rewards using metrics such as classification accuracy, recall, ranking correlation, and separation metrics, and further explore the robustness of the inferred reward functions under various perturbations and noisy conditions.
@@
By exposing the hidden incentives that drive LLM behavior, our approach not only enhances our theoretical understanding of RLHF but also paves the way for practical interventions that can mitigate risks associated with model deployment. In the remainder of the paper, we detail the technical aspects of our IRL framework, present extensive empirical results, and discuss future directions for integrating IRL into the broader landscape of model auditing and safety research.
@@ 2 Methodology
Let $\mathcal{D}=\{(c_{i},y_{i})\}_{i=1}^{N}$ be a dataset where $c_{i}$ represents a comment and $y_{i}\in\{0,1\}$ denotes the label. In our work, we focus on the task of toxicity (0 = non-toxic, 1 = toxic). We construct a balanced dataset $\mathcal{D}_{\text{bal}}\subset\mathcal{D}$ containing equal numbers of toxic and non-toxic samples. Each data point is further split into prompt-output pairs $(p_{i},o_{i})$ for training and evaluation purposes. The resulting $\mathcal{D}_{\text{bal}}$ dataset forms the basis for both RLHF fine-tuning, IRL reward extraction and evaluation.
@@ 2.2 Extracting the Ground Truth Reward Model $R^{*}$ and Fine Tuning LLM using $R^{*}$
Extracting Ground Truth $R^{*}$. An effective reward function is fundamental to the RLHF methodology, acting as an automated substitute for human input. We assume the true reward function $R^{*}$ is unknown at training time, as is the case in most practical scenarios, but can be parameterised by the classifier $f_{\theta}:\mathcal{O}\rightarrow\mathbb{R}$, where $\mathcal{O}$ is the space of text outputs. That is, $R^{*}$ can distinguish between toxic and non-toxic comments (0 = non-toxic, 1 = toxic). Formally, the reward function is: $R^{*}(o)=-f_{\theta}(o)$, where $o\in\mathcal{O}$. This ensures that outputs classified as less toxic receive higher rewards during RLHF.
Fine-Tuning LLMs with RLHF Using $R^{*}$. Given an LLM $\pi_{\phi}$ we finetune the model using RLHF for toxicity reduction. The objective is to maximize the expected reward provided by $R^{*}$. Our custom reward function encourages the model to generate content less toxic than the original while maintaining relevance to the prompt. The training process involves iterative sampling of prompts, generating responses, and updating the model parameters to maximize expected rewards. Specifically, given a prompt $p\in\mathcal{P}$, the model samples an output $o\sim\pi_{\phi}(\cdot|p)$ and updates $\pi_{\phi}$ to maximize $\mathbb{E}_{o\sim\pi_{\phi}(\cdot|p)}[R^{*}(o)]$ via proximal policy optimization (PPO). This yields an RLHF policy $\pi_{\text{E}}$ that prefers non-toxic completions while retaining relevance to the prompt. While we expect $\pi_{\text{E}}$ to overfit due to dataset size and model scale, this controlled setting facilitates our primary goal of testing reward recoverability using IRL.
@@ 2.3 Inverse RL for Approximating Model Incentives in LLMs
Let $\pi_{E}$ denote the RLHF-trained LLM. We treat $\pi_{E}$ as the expert policy generating trajectories $\{\tau_{E}\}$ over an MDP with: States $s_{t}=(p,a_{1},\dots,a_{t})$: partial sequences of prompts and outputs; Actions $a_{t}\in\mathcal{V}$: token outputs; Transition dynamics via autoregressive sampling from $\pi_{E}$; and Reward $R^{*}(s)=w^{T}\phi(s)$, with $w\in\mathbb{R}^{d}$ and $\phi:\mathcal{S}\to\mathbb{R}^{d}$ extracting interpretable features. Let $\mu_{E}$ and $\mu(\pi)$ represent the expected feature counts for the expert (LLM) policy and generated policies, respectively.
Using Max-Margin IRL to Approximate $\hat{R}$. We aim to extract an approximation $\hat{R}$ of ground truth $R^{*}$ from the fine-tuned model $\pi_{E}$ using Max-Margin IRL. Given a set of paired samples $(o^{+},o^{-})$ where $o^{+}\sim\pi_{\text{E}}$ (non-toxic) and $o^{-}\sim\pi_{\text{base}}$ (toxic) for the same prompt $p$, we seek to learn $\hat{R}:\mathcal{O}\rightarrow\mathbb{R}$ such that, $\hat{R}(o^{+})-\hat{R}(o^{-})\geq\delta$ for some positive margin $\delta$. We parameterize $\hat{R}$ as a reward head on top of the base LLM encoder $\pi_{\text{base}}$, adding a linear transformation from the hidden representation $h(o)$ to a scalar reward, $\hat{R}(o)=w^{\top}h(o)+b$, where $w$ and $b$ are learnable parameters. We optimize $\hat{R}$ using an asymmetric max-margin loss:
@@
where $x=\hat{R}(o^{+})-\hat{R}(o^{-})$. The loss penalizes violations of the margin constraint more heavily when $\hat{R}$ assigns higher rewards to toxic outputs than to non-toxic ones, reinforcing the sensitivity of $\hat{R}$ to undesirable behaviors.
@@
At each iteration, $R_{t}$ evaluates toxic/non-toxic samples, minimizing an asymmetric loss favoring non-toxic classifications. Gradients are backpropagated through $\hat{R}$. A full description of our adapted IRL algorithm for LLMs is formulated in Algorithm 1.
Analysing the Features of the Inferred Reward. The features $\phi(s)$ encode interpretable state attributes, such as n-gram statistics, coherence, relevance, sentiment, and toxicity scores. The inferred reward $\hat{R}(s)=\hat{w}^{T}\phi(s)$ provides insight into the LLM’s learned incentives, with weights $w$ revealing influences on its outputs, biases, safety concerns, or decision-making objectives. Analyzing $\hat{w}$ may uncover reward hacking, reliance on shallow heuristics, or spurious features. In what follows, we first present a simplified toy demonstration, followed by large-scale experiments to explore these issues.
@@ 3.2 Fine-Tuning and Baseline Generation
Fine-Tuning Procedure. We fine-tuned the Pythia-70Mn [^2] Link: lomahony/eleuther-pythia70m-hh-sft causal language model to generate both toxic and non-toxic adjective completions. Each prompt appeared twice: once with a toxic adjective and once with a non-toxic one. This setup tests whether the model can learn to differentiate between helpful and harmful completions, establishing a preference structure between toxic and non-toxic outputs. This structure is crucial for learning a reward function $R$ that favors non-toxic completions, similar to the behavior in RLHF fine-tuned models.
@@
[…]
@@ 3.4 Results of Demonstration
Precision 0.73
Recall 0.92
Kendall Tau 0.39
@@
IRL learns a reward model that generalizes well. As shown in Table 1, the IRL reward model achieved a precision of 0.733974. This is particularly notable given the relatively small number of unique toxic adjectives generated by the baseline model (Table 3), indicating that the reward model was capable of generalizing beyond specific training instances. Additionally, the IRL reward model achieves a recall of over 0.916000 and a Kendall Tau rank correlation of 0.396553, indicating its ability to maintain correct ordering between samples according to reward value. The violin and scatter plots in Figure 1 show clear separation between toxic and non-toxic categories, with the reward function assigning higher values to desirable (non-toxic) outputs.
Figure 1: Evaluation of the IRL: extracted reward function on toxic and non-toxic adjective completions. Left: Scatter plot comparing ground truth rewards (x-axis) to IRL-extracted rewards (y-axis), revealing strong alignment and effective ranking. Middle: Violin plot showing clear separation between toxic and non-toxic samples in terms of extracted reward. Right: Confusion matrix indicating strong classification performance (Precision: 0.73, Recall: 0.91), with high accuracy in identifying toxic outputs and some false negatives among non-toxic samples.
@@
IRL is robust to out-of-distribution data and can extract meaningful rewards even in imbalanced data settings. We assess IRL’s robustness by varying the toxic-to-non-toxic ratio in training. In Figure 2(a, b, c), we fix non-toxic samples and incrementally add toxic ones. As toxicity increases, performance improves: (i) Precision peaks at  77%, (ii) Recall remains consistently high, and (iii) Kendall Tau steadily rises, showing the reward model becomes better at capturing the relative ordering of toxic vs. non-toxic samples as it gains more contrastive signal. Even with no toxic data (toxic fraction = 0), the model achieves 0̃.64 precision. This demonstrates that the model is able to infer a meaningful reward structure from only non-toxic samples, effectively learning what is “good” without needing explicit negative examples. The IRL reward model is able to distinguish between toxic samples which are outside of distribution. As more toxic samples are incrementally added, performance improves.
@@ 4 Experiments
Language Models. The experiments use two Pythia language models (70M and 410M parameters) that underwent one epoch of Supervised Fine-Tuning (SFT) on the Anthropic Helpful and Harmless (HH) dataset. These models, designed for interpretability research, share standardized training and data for reproducibility. Starting with SFT models ensures they generate helpful and safe content without reinforcement learning, mirroring real-world RLHF applications. Using two model sizes allows analysis of how scale affects toxicity reduction and IRL reward learning.
Experimental Setup. We implement RLHF using the TRLx library, adapting it for toxicity reduction. The true reward function $R^{*}$ is implemented to encourage the model to generate less toxic content while maintaining relevance. We use PPO for training, with a cosine learning rate schedule and AdamW optimizer. A KL divergence term is incorporated to prevent extreme policy shifts. Key metrics such as returns/mean and reward/mean are monitored throughout training to assess toxicity reduction and output quality. Table 5 details the hyperparameters used for fine-tuning the 70M and 410M models, they are tuned to balance exploration and exploitation.

### v3

@@ Abstract
Large language models (LLMs) trained with Reinforcement Learning from Human Feedback (RLHF) have demonstrated remarkable capabilities, but their underlying reward functions and decision-making processes remain opaque. This paper introduces a novel approach to interpreting LLMs by applying inverse reinforcement learning (IRL) to recover their implicit reward functions. We conduct experiments on toxicity-aligned LLMs of varying sizes, extracting reward models that achieve up to 85% accuracy in predicting human preferences. Our analysis reveals key insights into the non-identifiability of reward functions, the relationship between model size and interpretability, and potential pitfalls in the RLHF process. We demonstrate that IRL-derived reward models can be used to fine-tune new LLMs, resulting in comparable or improved performance on toxicity benchmarks. This work provides a new lens for understanding and improving LLM alignment, with implications for the responsible development and deployment of these powerful systems. [^2] Code for our paper can be found at https://github.com/ai4ai-lab/irl_for_llms
@@ 1 Introduction
Inverse Reinforcement Learning (IRL) is widely used in robotics and control theory to infer latent reward structures from observed behavior (Ng & Russell, 2000). By treating observed actions as expert demonstrations, IRL techniques aim to reverse-engineer the hidden reward function that an agent is implicitly optimizing. In this work, we propose a novel application of IRL to LLMs trained via RLHF. We posit that if the outputs of an RLHF-trained LLM can be interpreted as demonstrations from an ”expert” policy, then IRL methods—particularly those based on maximum margin formulations—can be employed to recover the hidden reward functions that guided the training process. This approach provides insights into decision-making processes and enhances model auditing.
@@
We employ Maximum Margin IRL to learn a reward model $\hat{R}:\mathcal{O}\rightarrow\mathbb{R}$ that distinguishes between preferred (non-toxic) and suboptimal (toxic) responses. By parameterizing $\hat{R}$ using hidden representations from a base LLM and optimizing it with an asymmetric margin loss, we enforce clear reward separation while addressing non-identifiability challenges. Our experiments focus on toxicity reduction, evaluating Pythia models (70M and 410M parameters) (Biderman et al., 2023) on Jigsaw Toxicity and RealToxicityPrompts benchmarks. Results show that IRL-extracted reward functions closely align with human judgments. We quantitatively evaluate the recovered rewards using metrics such as classification accuracy, recall, ranking correlation, and separation metrics, and further explore the robustness of the inferred reward functions under various perturbations and noisy conditions.
@@
By exposing the hidden incentives that drive LLM behavior, our approach not only enhances our theoretical understanding of RLHF but also paves the way for practical interventions that can mitigate model deployment risk. The rest of the paper details the technical aspects of our IRL framework, present extensive empirical results, and discusses future directions for integrating IRL into the broader landscape of model auditing and safety research.
@@ 2 Methodology
Let $\mathcal{D}=\{(c_{i},y_{i})\}_{i=1}^{N}$ be a dataset where $c_{i}$ represents a comment and $y_{i}\in\mathbb{R}$ denotes the label. In our work, we focus on the task of toxicity (lower values correspond to non-toxic, while higher values to toxic). We construct a balanced dataset $\mathcal{D}_{\text{bal}}\subset\mathcal{D}$ containing equal numbers of toxic and non-toxic samples. Each data point is further split into prompt-output pairs $(p_{i},o_{i})$ for training and evaluation purposes. The resulting $\mathcal{D}_{\text{bal}}$ dataset forms the basis for both RLHF fine-tuning, IRL reward extraction and evaluation.
@@ 2.2 Extracting the Ground Truth Reward Model $R^{*}$ and Fine Tuning LLM using $R^{*}$
Extracting Ground Truth $R^{*}$. An effective reward function is fundamental to the RLHF methodology, acting as an automated substitute for human input. We assume the true reward function $R^{*}$ is unknown at training time, as is the case in most practical scenarios, but can be parameterised by the function $f_{\theta}:\mathcal{O}\rightarrow\mathbb{R}$, where $\mathcal{O}$ is the space of text outputs. That is, $R^{*}$ can distinguish between toxic and non-toxic comments (0 = non-toxic, 1 = toxic). Formally, the reward function is: $R^{*}(o)=-f_{\theta}(o)$, where $o\in\mathcal{O}$. This ensures that outputs classified as less toxic receive higher rewards during RLHF.
Fine-Tuning LLMs with RLHF Using $R^{*}$. Given an LLM $\pi_{\phi}$ we finetune the model using RLHF for toxicity reduction. The objective is to maximize the expected reward provided by $R^{*}$. Our custom reward function encourages the model to generate content less toxic than the original while maintaining relevance to the prompt. The training process involves iterative sampling of prompts, generating responses, and updating the model parameters to maximize expected rewards. Specifically, given a prompt $p\in\mathcal{P}$, the model samples an output $o\sim\pi_{\phi}(\cdot|p)$ and updates $\pi_{\phi}$ to maximize $\mathbb{E}_{o\sim\pi_{\phi}(\cdot|p)}[R^{*}(o)]$ via proximal policy optimization (PPO) (Schulman et al., 2017). This yields an RLHF policy $\pi_{\text{E}}$ that prefers non-toxic completions while retaining relevance to the prompt. While we expect $\pi_{\text{E}}$ to overfit due to dataset size and model scale, this controlled setting facilitates our primary goal of testing reward recoverability using IRL.
@@ 2.3 Inverse RL for Approximating Model Incentives in LLMs
Let $\pi_{E}$ denote the RLHF-trained LLM. We treat $\pi_{E}$ as the expert policy. The outputs generated by this policy can be considered expert trajectories $\{\tau_{E}\}$ over an MDP with: States $s_{t}=(p,a_{1},\dots,a_{t})$: partial sequences of prompts and outputs; Actions $a_{t}\in\mathcal{V}$: token outputs; Transition dynamics via autoregressive sampling from $\pi_{E}$; and Reward $R^{*}(s)=w^{T}\phi(s)$, with $w\in\mathbb{R}^{d}$ and $\phi:\mathcal{S}\to\mathbb{R}^{d}$ extracting interpretable features. Let $\mu_{E}$ and $\mu(\pi)$ represent the expected feature counts for the expert (LLM) policy and generated policies, respectively.
Using Max-Margin IRL to Approximate $\hat{R}$. We aim to extract an approximation $\hat{R}$ of ground truth $R^{*}$ from the fine-tuned model $\pi_{E}$ using Max-Margin IRL. Given a set of paired samples $(o^{+},o^{-})$ where $o^{+}\sim\pi_{\text{E}}$ (non-toxic) and $o^{-}\sim\pi_{\text{base}}$ (toxic) for the same prompt $p$, we seek to learn $\hat{R}:\mathcal{O}\rightarrow\mathbb{R}$ such that, $\hat{R}(o^{+})-\hat{R}(o^{-})\geq\delta$ for some positive margin $\delta$.
We parameterize $\hat{R}$ as a reward head on top of the base LLM encoder $\pi_{\text{base}}$. Specifically, for each output $o\in\mathcal{O}$, we extract a hidden representation $h(o)$ from the model, which serves as our feature function $\phi(s)$ in the IRL formulation. A linear layer maps this embedding to a scalar reward: $\hat{R}(o)=w^{\top}h(o)+b$, where $w$ and $b$ are trainable parameters. The expert feature expectations $\mu_{E}$ in Algorithm 1 are therefore the mean of these hidden representations $h(o)$ across all expert-generated trajectories.
We optimize $\hat{R}$ using an asymmetric max-margin loss (Shah et al., 2022):
@@
where $x=\hat{R}(o^{+})-\hat{R}(o^{-})$. The loss function is asymmetric, penalizing violations of the margin constraint more heavily when the reward model $\hat{R}$ assigns higher rewards to toxic outputs than to non-toxic ones. This design encourages $\hat{R}$ to be particularly sensitive to undesirable behaviors, reflecting the safety-critical nature of the alignment task. Inspired by max-margin IRL (Ratliff et al., 2006), the loss imposes a steeper penalty when x<0x<0—i.e., when a toxic output is incorrectly preferred—thereby pushing the model to learn nuanced preference gradients rather than relying on coarse, discrete class labels. At each iteration, $R_{t}$ evaluates toxic/non-toxic samples, minimizing an asymmetric loss favoring non-toxic classifications. Gradients are backpropagated through $\hat{R}$. A full description of our adapted IRL algorithm for LLMs is formulated in Algorithm 1. Note that we allow the user to set the convergence threshold $\epsilon$, since empirical performance is typically governed more by the informativeness of the expert trajectories than by their number.
@@
Analysing the Features of the Inferred Reward. The features $\phi(s)$ are task-specific and designed to capture key state attributes. The reward model is trained to minimize the difference in feature expectations (under $\phi$) between expert and generated trajectories. $\phi(s)$ can encode interpretable properties like n-gram statistics, coherence, relevance, sentiment, or toxicity. For alignment tasks such as factuality, it might capture slur indicators, coherence, or domain cues. In our case, we use token embeddings as $\phi$, since subword-level representations capture rich semantic and stylistic signals (e.g., toxicity) and are easy to implement. Typically, $\phi$ is task-dependent and chosen by the practitioner. The choice of a linear reward model $\hat{R}(s)=\hat{w}^{T}\phi(s)$ is deliberate: each coordinate of $\phi$ has a clear, human-interpretable meaning. The learned weights $\hat{w}$ assign positive or negative valence to features, revealing what the LLM promotes or suppresses—shedding light on its biases, safety concerns, and objectives. Analyzing $\hat{w}$ can uncover reward hacking, shallow heuristics, or spurious correlations. In what follows, we first present a simplified toy demonstration, followed by large-scale experiments to explore these issues.
@@ 3.2 Fine-Tuning and Baseline Generation
Fine-Tuning Procedure. We fine-tuned the Pythia-70Mn [^3] Link: lomahony/eleuther-pythia70m-hh-sft causal language model to generate both toxic and non-toxic adjective completions. Each prompt appeared twice: once with a toxic adjective and once with a non-toxic one. This setup tests whether the model can learn to differentiate between helpful and harmful completions, establishing a preference structure between toxic and non-toxic outputs. This structure is crucial for learning a reward function $\hat{R}$ that favors non-toxic completions, similar to the behavior in RLHF fine-tuned models.
@@
RLHF and Model Training. To enforce non-toxic language, we fine-tune a pretrained model using RLHF, encouraging non-toxic and discouraging toxic adjectives. Training runs for 1 epoch with Adam (learning rate $1\mathrm{e}{-6}$), $\text{top\_p}=0.95$, $\text{temperature}=1.4$, and entropy regularization ($\lambda=0.005$). The dataset ($80/20$ split) is shuffled before splitting, with RLHF updates applied only to training data, while validation monitors average reward to prevent overfitting. The reward function assigns positive scores to non-toxic adjectives, negative to toxic ones, and intermediate penalties when both appear. This shifts model outputs from toxic (e.g., horrible) to socially acceptable alternatives (e.g., productive, pleasant), embedding a reward signal favoring non-toxic language. Table 4 in Appendix C summarizes the adjective generation behavior of the RLHF-trained model on the 250 prompts used during training.
@@ 3.4 Results of Demonstration
Precision 0.75
Recall 0.90
Kendall Tau 0.40
@@
IRL learns a reward model that generalizes well. As shown in Table 1, the IRL reward model achieved a precision of 0.745033. This is particularly notable given the relatively small number of unique toxic adjectives generated by the baseline model (Table 3), indicating that the reward model was capable of generalizing beyond specific training instances. Additionally, the IRL reward model achieves a recall of over 0.900000 and a Kendall Tau rank correlation of 0.401972, indicating its ability to maintain correct ordering between samples according to reward value. The violin and scatter plots in Figure 1 show clear separation between toxic and non-toxic categories, with the reward function assigning higher values to desirable (non-toxic) outputs.
Figure 1: Evaluation of the IRL: extracted reward function on toxic and non-toxic adjective completions. Left: Scatter plot comparing ground truth rewards (x-axis) to IRL-extracted rewards (y-axis), revealing strong alignment and effective ranking. Middle: Violin plot showing clear separation between toxic and non-toxic samples in terms of extracted reward. Right: Confusion matrix indicating strong classification performance (Precision: 0.75, Recall: 0.90), with high accuracy in identifying toxic outputs and some false negatives among non-toxic samples.
@@
IRL is robust to out-of-distribution data and can extract meaningful rewards even in imbalanced data settings. We assess IRL’s robustness by varying the toxic-to-non-toxic ratio in training. In Figure 2(a, b, c), we fix non-toxic samples and incrementally add toxic ones. As toxicity increases, performance improves: (i) Precision peaks at  74%, (ii) Recall remains consistently high, and (iii) Kendall Tau steadily rises, showing the reward model becomes better at capturing the relative ordering of toxic vs. non-toxic samples as it gains more contrastive signal. Even with no toxic data (toxic fraction = 0), the model achieves 0̃.64 precision. This demonstrates that the model is able to infer a meaningful reward structure from only non-toxic samples, effectively learning what is “good” without needing explicit negative examples. The IRL reward model is able to distinguish between toxic samples which are outside of distribution. As more toxic samples are incrementally added, performance improves.
@@
4 Experiments
For our experiments, we focus on toxicity reduction since it is one of the most fundamental alignment objectives addressed via RLHF and serves as a primary benchmark in both industrial and academic evaluations of safety-aligned LLMs (Ouyang et al., 2022; Wang et al., 2023). That said, our method is agnostic to the specific alignment signal and toxicity signals could without loss of generality be replaced by signals for bias or factuality features (e.g., demographic association signals, named-entity consistency metrics).
Language Models. The experiments use two Pythia language models (70M and 410M parameters) (Biderman et al., 2023) that underwent one epoch of Supervised Fine-Tuning (SFT) (Ouyang et al., 2022) on the Anthropic Helpful and Harmless (HH) dataset (Bai et al., 2022a). These models, designed for interpretability research, share standardized training and data for reproducibility. Starting with SFT models ensures they generate helpful and safe content without reinforcement learning, mirroring real-world RLHF applications. Using two model sizes allows analysis of how scale affects toxicity reduction and IRL reward learning. For test-time, we use Jigsaw-Toxicity and RealToxicityPrompts as they are public, orthogonal to Anthropic-HH, and benchmark the exact alignment dimension (toxicity) under consideration. For an unknown closed-source model the practitioner should select any prompt corpus that captures the property of interest (e.g., factuality, helpfulness).
Experimental Setup. We implement RLHF using the TRLx library, adapting it for toxicity reduction. The true reward function $R^{*}$ encourages the model to generate less toxic yet relevant outputs. We use PPO (Schulman et al., 2017) for training, with a cosine learning rate schedule and AdamW optimizer. A KL divergence term is incorporated to prevent extreme policy shifts. Key metrics (returns/mean and reward/mean) are monitored throughout training to assess toxicity reduction and output quality. Our IRL formulation makes two assumptions that are unavoidably required for any reward-recovery method: (i) Access to expert trajectories emitted by the RLHF-tuned policy $\pi_{E}$ and (ii) The feature map $\phi$ that is rich enough to separate desirable from undesirable behaviors. Neither assumption however ties the method to a known pre-training corpus or alignment dataset.